2x+2y+3z=13
-3x+4y-z=5
5x-3y+z=2
solve a system of 3 linear equations in 3 variables by the addition method. I keep coming up with fractions for x variable. i must be missing something in one of steps.

Answer by scott8148(5891)

What's wrong with fractions?

adding (equation 2) to (equation 3) gives 2x+y=7 (a)

multiplying (equation 2) by 3 and adding it to (equation 1) gives -7x+14y=28, dividing this by 7 gives -x+2y=4 (b)

multiplying (equation (b)) by 2 and adding (equation (a)) gives 5y=15

so y=3,, x=2,, and z=1